OpenStudy (anonymous):

(1-3^-1)(-3^-1) Steps?

OpenStudy (anonymous):

(1-1/3)*(-1/3) = 2/3*(-1/3)=-2/9

OpenStudy (anonymous):

Oh, I see how that works. Thank you very much.